### v1

# GeoGen: A Two-stage Coarse-to-Fine Framework for Fine-grained Synthetic Location-based Social Network Trajectory Generation

Abstract
Location-Based Social Network (LBSN) check-in trajectory data are important for many practical applications like POI recommendation, advertising, and pandemic intervention. However, the high collection costs and ever-increasing privacy concerns prevent us from accessing large-scale LBSN trajectory data. The recent advances in synthetic data generation provide us with a new opportunity to achieve this, which utilizes generative AI to generate synthetic data that preserves the characteristics of real data while ensuring privacy protection. However, generating synthetic LBSN check-in trajectories remains challenging due to their spatially discrete, temporally irregular nature and the complex spatio-temporal patterns caused by sparse activities and uncertain human mobility. To address this challenge, we propose GeoGen, a two-stage coarse-to-fine framework for large-scale LBSN check-in trajectory generation. In the first stage, we reconstruct spatially continuous, temporally regular latent movement sequences from the original LBSN check-in trajectories and then design a Sparsity-aware Spatio-temporal Diffusion model (S2TDiff) with an efficient denosing network to learn their underlying behavioral patterns. In the second stage, we design Coarse2FineNet, a Transformer-based Seq2Seq architecture equipped with a dynamic context fusion mechanism in the encoder and a multi-task hybrid-head decoder, which generates fine-grained LBSN trajectories based on coarse-grained latent movement sequences by modeling semantic relevance and behavioral uncertainty. Extensive experiments on four real-world datasets show that GeoGen excels state-of-the-art models for both fidelity and utility evaluation, e.g., it increases over 69% and 55% in distance and radius metrics on the FS-TKY dataset.
Introduction
Fine-grained location-based social network (LBSN) check-in trajectory data are essential for many real-world applications such as next Point of Interest (POI) recommendation (Jiang et al. 2024), urban mobility understanding (Wang et al. 2019), business location selection (Xie et al. 2016), advertising (Jeon et al. 2021), and pandemic intervention (Hao et al. 2020). However, accessing large-scale LBSN data has become increasingly challenging due to high collection/purchasing costs and growing privacy concerns. Therefore, to support large-scale LBSN data access, it is crucial to develop methods that can generate synthetic LBSN check-in trajectories that maintain high fidelity and utility while safeguarding user privacy.
To date, some models have been proposed (Isaacman et al. 2012; Jiang et al. 2016; Yin et al. 2017; Ouyang et al. 2018; Zhu et al. 2023b) to generate LBSN check-in trajectory data. For instance, Yin et al. (Yin et al. 2017) and Jiang et al. (Jiang et al. 2016) employed physics-based models like hidden Markov models to synthesize activity trajectories. Generative adversarial networks (GANs) (Ouyang et al. 2018; Feng et al. 2020; Yu et al. 2017; Yuan et al. 2022) have demonstrated substantial advances in generating LBSN check-in trajectories. However, these approaches inevitably suffer from an unstable and extremely long training procedure. The impressive achievements of diffusion models in various tasks, such as image (Dhariwal and Nichol 2021; Kulikov et al. 2023) and audio (Huang et al. 2022) generation, have sparked interest in their applications to trajectory generation. Several diffusion-based models (Zhu et al. 2023b, 2024; Liu et al. 2023) have been proposed to generate spatially continuous, temporally regular GPS trajectories with fixed intervals between consecutive data points and fixed lengths (e.g., 100 data points per trajectory). These models have demonstrated strong performance, highlighting the great potential of diffusion models for generating LBSN check-in trajectories.
Figure 1: Example of a discrete LBSN trajectory with irregular intervals (blue); and a continuous GPS trajectory with a fixed interval of 30 seconds (red).
However, it is nontrivial to adapt diffusion models to LBSN check-in trajectory data generation due to its two unique natures: (i) Temporally irregular: As shown in Figure 1, in contrast to regularly sampled GPS trajectories, LBSN check-in trajectories exhibit complex spatio-temporal patterns due to uncertain human mobility and check-in behavior, leading to variable time intervals between points and unfixed trajectory lengths. (ii) Spatially discrete: Diffusion models are originally designed to generate fixed-length data in the continuous domain, while POIs are from a discrete spatial space, so it is challenging to adapt existing diffusion models designed for regularly-sampled trajectory generation (Zhu et al. 2023b, 2024; Liu et al. 2023) to LBSN check-in trajectories. In addition, a large time interval setting (e.g., 60 minutes) may decrease the accuracy for LBSN check-in trajectory generation, while a small interval (e.g., 1 minute) may significantly decrease the generation efficiency.
To address these challenges, we propose GeoGen, a two-stage, coarse-to-fine framework for synthetic LBSN trajectory data generation. In the first stage, to adapt diffusion models to irregularly-sampled LBSN trajectory data, we reconstruct spatially continuous and temporally regular latent movement sequences from LBSN check-in trajectories through interpolation with a pre-determined coarse-grained interval. We then design a Sparsity-aware Spatio-temporal Diffusion model (S2TDiff), which includes a novel and efficient Spatially-Aware Sparsely-Gated U-Net (SASG-UNet) as its denoising network. SASG-UNet replaces standard convolutions with hierarchical 1D blocks to capture multi-scale patterns and integrates a specialized S2G Attention module to address the unique characteristics of the latent movement sequences. In the second stage, we propose a Transformer-based Seq2Seq architecture Coarse2FineNet that auto-regressively generates fine-grained check-in points based on the coarse-grained latent movement sequences. Coarse2FineNet incorporates a POI context-aware encoder that captures rich contextual representations of the latent movement sequence, and a Multi-task Hybrid-Head Decoder that generates the next POI and its fine-grained timestamp. In the encoder, we introduce a dynamic context fusion mechanism that aligns the input latent movement sequence with POIs exhibiting diverse spatio-temporal characteristics, integrating them into unified contextual representations. In the decoder, instead of directly predicting the next check-in point, we leverage a temporal point process to model the user behavior uncertainty underlying LBSN check-in trajectories. This two-stage approach effectively resolves a key trade-off between computational efficiency and data quality, enabling the generation of large-scale synthetic LBSN check-in trajectories.
The key contributions of this paper are as follows:
  • Conceptually, we focus on developing innovative diffusion models to efficiently generate large-scale fine-grained LBSN check-in trajectories, which are inherently spatially discrete and temporally irregular.
  • Technically, we propose a two-stage coarse-to-fine framework called GeoGen for synthetic LBSN trajectory generation. In the first stage, we design a novel S2TDiff model with an efficient SASG-UNet as a denoising network to generate coarse-grained latent movement sequences. In the second stage, we design a Transformer-based Seq2Seq architecture, Coarse2FineNet, which includes a POI context-aware encoder and a multi-task hybrid-head decoder to generate fine-grained POI visits with precise, non-uniform timestamps.
  • Experimentally, we conduct extensive experiments on four real-world LBSN check-in datasets. The results demonstrate that GeoGen significantly outperforms state-of-the-art models in data fidelity and show high utility for downstream tasks like next location prediction.
PRELIMINARIES
Problem Statement
Definition 1. (LBSN Check-in Trajectories). An LBSN check-in trajectory is described as a POI visit sequence $s=[x_{1},x_{2},\ldots]$ of a user. The $i$th element $x_{i}$ is a check-in point defined as a tuple $(p_{i},t_{i})$, where $p_{i}$ denotes a unique POI location ID corresponding to a geographic coordinate $g_{p_{i}}$, and $t_{i}$ is the check-in timestamp.
Definition 2. (Trajectory Granularity) The granularity of a trajectory is its smallest recorded time unit. Coarse-grained trajectories use large time units (e.g., hours), while fine-grained trajectories use smaller ones (e.g., minutes). This work aims to generate fine-grained LBSN check-in trajectories with a minute-level granularity.
Definition 3. (LBSN Check-in Trajectory Generation). Given a real-world dataset containing $m$ fine-grained LBSN check-in trajectories, denoted as $S=[s_{1},s_{2},\ldots]$, the LBSN check-in trajectory generation task aims to generate a synthetic trajectory dataset $\hat{S}=[\hat{s}_{1},\hat{s}_{2},\ldots]$ that maintains the spatio-temporal characteristics and distributions of $S$, as well as high data utility for downstream applications.
Figure 2: The overall architecture of our proposed GeoGen framework consists of two main stages. In the first stage, within S2TDiff, each reconstructed latent movement sequence $s_{\text{lat}}$ is first transformed into pure noise $s_{\text{lat}}^{N}$. During inference, SASG-UNet progressively denoises a sampled noise sequence to generate a synthetic latent trajectory. In the second stage, the context-aware encoder of Coarse2FineNet extracts contextualized representations from the generated latent movement sequence. The decoder then auto-regressively generates each POI and fine-grained timestamp pair, conditioned on both the contextual representations and the partially generated LBSN check-in trajectory. FFN and TPP refer to the feed-forward network and temporal point process, respectively.
Denoising Diffusion Probabilistic Model
Denoising Diffusion Probabilistic Model (Ho, Jain, and Abbeel 2020) encompass two processes: (i) a forward diffusion process that gradually introduces noise into the data, and (ii) a backward diffusion process that reconstructs the original data from its noisy state. The forward diffusion process is a non-trainable Markov chain that transforms real sample $\mathbf{x}^{0}$ into latent variables $\mathbf{x}^{1},\ldots,\mathbf{x}^{N}$, which can be represented as:
$q(\mathbf{x}^{n}|\mathbf{x}^{n-1})=\mathcal{N}(\mathbf{x}^{n};\sqrt{1-\beta_{n}}\mathbf{x}^{n-1},\beta_{n}\mathbf{I}),$ (1)
where $\beta_{1},\ldots,\beta_{N}\in(0,1)$ is a pre-determined variance schedule. $\mathbf{x}^{n}$ can be directly sampled from $\mathbf{x}^{0}$ with $q(\mathbf{x}^{n}|\mathbf{x}^{0})=\mathcal{N}(\mathbf{x}^{n};\sqrt{\overline{\alpha}_{n}}\mathbf{x}^{0},(1-\overline{\alpha}_{n})\mathbf{I})$, where $\alpha_{n}=1-\beta_{n}$ and $\overline{\alpha}_{n}=\prod_{n=1}^{N}\alpha_{n}$. With the application of the reparameterization, $\mathbf{x}^{n}$ can be represented as $\mathbf{x}^{n}=\sqrt{\overline{\alpha}_{n}}\mathbf{x}^{0}+\sqrt{1-\overline{\alpha}_{n}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. The backward diffusion process is a trainable Markov chain that aims to recover $\mathbf{x}^{0}$ from $\mathbf{x}^{N}$, which can be formulated as:
$p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})=\mathcal{N}(\mathbf{x}^{n-1};\mu_{\theta}(\mathbf{x}^{n},n),\sigma_{\theta}(\mathbf{x}^{n},n)\mathbf{I}),$ (2)
where $\mu_{\theta}(\mathbf{x}^{n},n)$ and $\sigma_{\theta}(\mathbf{x}^{n},n)$ are the mean and variance predicted normally by a neural network parameterized by $\theta$. The loss function is formalized as:
$\mathbb{E}_{\mathbf{x}^{0},\boldsymbol{\epsilon}}\left[\frac{\beta_{n}^{2}}{2\Sigma_{\theta}\alpha_{n}(1-\overline{\alpha}_{n})}\left\|\boldsymbol{\epsilon}-\epsilon_{\theta}\left(\sqrt{\overline{\alpha}_{n}}\mathbf{x}^{0}+\sqrt{1-\overline{\alpha}_{n}}\boldsymbol{\epsilon},n\right)\right\|^{2}\right],$ (3)
where $\epsilon_{\theta}$ is a neural network for predicting sampled $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. After trained, trajectory generation is conducted by progressively sampling $\mathbf{x}^{n-1}$ from distribution $p_{\theta}(\mathbf{x}^{n-1}|\mathbf{x}^{n})$ until reach $\mathbf{x}^{0}$ by computing:
$\mathbf{x}^{n-1}=\frac{1}{\sqrt{\alpha_{n}}}\left(\mathbf{x}^{n}-\frac{\beta_{n}}{\sqrt{1-\alpha_{n}}}\epsilon_{\theta}(\mathbf{x}^{n},n)\right)+\sqrt{\Sigma_{\theta}}\mathbf{z},$ (4)
where $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ for $n\in[2,N]$, and $\mathbf{z}=\mathbf{0}$ when $n=1$.
Methodology
In this part, we introduce the detailed design of GeoGen. An overall pipeline of GeoGen is shown in Figure 3.
Coarse-grained Latent Movement Sequence Generation
To adapt diffusion models to irregularly-sampled LBSN trajectory data, we first convert original fine-grained LBSN trajectories into coarse-grained spatially continuous and temporally regular latent movement sequences. We then design a Sparsity-aware Spatio-temporal Diffusion model (S2TDiff) to generate coarse-grained latent movement sequences by learning underlying spatio-temporal distributions and physical constraints.
Coarse-grained Latent Movement Sequence Reconstruction
We first reconstruct each real LBSN check-in trajectory $s$ into a latent movement sequence, $s_{\mathit{lat}}=[(l_{1},c_{1}),\ldots,(l_{L},c_{L})]$. Given a time interval $\mathit{I}$, we divide the total duration into $L$ slots, where $L=[D/I]$, $D$ is the duration. For each slot $i$, the latent coordinate $l_{i}$ is the geographic mean of all check-ins $P_{i}$ falling within that time window ($l_{i}=\frac{1}{|P_{i}|}\sum_{(p_{j},t_{j})\in P_{i}}g_{p_{j}}$), and the intensity $c_{i}$ is their count ($c_{i}=|P_{i}|$). Next, we adopt linear interpolation between the nearest preceding ($i_{p}$) and subsequent ($i_{n}$) time slots with no check-ins ($c_{i}=0$), where the missing location is calculated as $l_{i}=l_{i_{p}}+((i-i_{p})/(i_{n}-i_{p}))\cdot(l_{i_{n}}-l_{i_{p}})$. For empty slots at the sequence boundaries, we apply circular interpolation, treating the trajectory as a loop to infer locations between the first and last observed points. This reformulation transforms sparse event data into a spatially continuous and temporally regular format, enabling compatibility with diffusion models while reducing memory and computational costs via temporal compression.
Sparsity-aware Spatio-temporal Diffusion Model
Figure 3: The architecture of the proposed Spatially-Aware Sparsely-Gated U-Net (SASG-UNet). The kernel sizes for the average pooling layers in the Hierarchical 1D Convolution layers are indicated in parentheses.
To capture the underlying spatio-temporal patterns, we propose a Sparsity-aware Spatio-temporal Diffusion model (S2TDiff), trained using latent movement sequences $S_{\text{lat}}=[s_{\text{lat},1},s_{\text{lat},2},\dots]$ as input. While interpolation yields spatially continuous sequences, it poses challenges of (1) diverse movement rates and (2) sparse intensity.
To address these challenges, we design a novel and efficient Spatially-Aware Sparsely-Gated U-Net (SASG-UNet) as the denoising backbone in S2TDiff. S2TDiff employs a U-Net backbone with hierarchical 1D convolution blocks that replace standard 2D convolutions. These blocks efficiently capture multi-resolution features using parallel branches of multi-scale average pooling and fixed-size 1D convolutions. At training/inference step $n$, the denoising network SASG-UNet $\epsilon_{\theta}$ takes a noised latent movement sequence as input and outputs the predicted noise: $\epsilon_{\theta}(s_{\text{lat}}^{n},n)$. To further tackle sparse intensity and diverse movement rates, we introduce a novel S2G Attention module within SASG-UNet’s skip connections, consisting of two complementary pathways: the Intensity-Gated Pathway and the Spatially-Aware Attention Pathway. Firstly, the Intensity-Gated Pathway handles sparse signals by with a lightweight convolutional sub-network to generate a dynamic gating signal $g=\sigma(\text{Conv1D}(\text{ReLU}(\text{Conv1D}(h_{\text{enc}}))))$, where $h_{\text{enc}}$ denotes encoder features and $\sigma$ is a sigmoid function. This gate adaptively modulates the features, allowing the network to prioritize learning from informative, sparse, non-zero intensity values. Secondly, the Spatially-Aware Attention Pathway accounts for diverse movement rates via a Local Spatial Bias $b$, where at index $i$, $b_{i}=\mathcal{F}_{\text{emb}}(d_{\text{H}}((s^{n}_{\text{lat},i},s^{n}_{\text{lat},i-1}))$, and $d_{\text{H}}$ is the Haversine distance (Maria et al. 2020) between consecutive points $s^{n}_{\text{lat},i}$, $s^{n}_{\text{lat},i-1}$ in the noisy input and $\mathcal{F}_{\text{emb}}$ is a learnable feed-forward network. Next, we add an attention mechanism, yielding the final refined features: $h_{\text{att}}=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$, where $Q$ and $K$ are the self-attention output features of the Intensity-Gated Pathway and $V$ is the bias $b$. This mechanism provides the model with a direct and efficient understanding of real-world movement dynamics, enabling it to handle sparse, irregular mobility patterns effectively.
Coarse2FineNet for Fine-grained LBSN Trajectory Generation with Context Awareness
In the second stage, we propose Coarse2FineNet, a Transformer-based Seq2Seq framework that translates coarse-grained latent movement sequences generated by S2TDiff into fine-grained LBSN check-in trajectories by jointly modeling the spatio-temporal features of the latent movement sequence and characteristics of POIs. Coarse2FineNet includes two key components: a Context-Aware Encoder and a Multi-task Hybrid-Head Decoder. The encoder captures a rich, contextualized representation of user mobility from coarse-grained latent movement sequences, while the decoder autoregressively generates a sequence of POI visits with fine-grained timestamps.
Context-Aware Encoder
To reduce complexity in the diffusion process, we compress the latent sequence $s_{\text{latent}}=\{(l_{i},c_{i})\}$ into $s_{\text{latent}}^{\prime}=\{(l^{\prime}_{i},t^{\prime}_{i})\mid c_{i}\geq\gamma\}$ by filtering points, where $l^{\prime}_{j}=l_{i_{j}}$ and $t^{\prime}_{j}=t_{i_{j}}=(i_{j}-0.5)\cdot\mathit{Int}$, $\gamma$ denotes intensity threshold, and $t_{i}$ denotes the midpoint time of slot $i$. We first obtain spatial embeddings $l_{vec}=\{Linear(l^{\prime}_{i})\}$ and temporal embedding $t_{vec}=\{T2V(t^{\prime}_{i})\}$, where $T2V$ denotes the Time2Vec method  (Kazemi et al. 2019) that captures both linear and cyclical temporal patterns.
To establish local POI context awareness, we design a Dynamic Context Fusion (DCF) module within the encoder. In DCF, a shared POI embedding codebook is built to encode multiple attributes, including geographic coordinates $p_{j}$, visit frequency vectors $f_{j}$, and category one-hot vector $c_{j}$ for POI $j$: $e_{\text{POI}}^{(j)}=W_{\text{latlon}}p_{j}+W_{\text{freq}}f_{j}+W_{\text{cat}}c_{j}$, where $W_{\text{latlon}},W_{\text{freq}},W_{\text{cat}}$ are learnable parameters. Then, we employ a dual attention mechanism to model spatial and temporal relationships separately. The spatial attention maps each coordinate $l^{\prime}_{i}$ from the filtered sequence to a distribution over POIs based on proximity. The attention weight $\alpha_{j}$ for each POI $j$ and coordinate $l_{i}$ is calculated as:
$\alpha_{i,j}^{(s)}=\frac{\exp(-\|l^{\prime}_{i}-g_{p_{j}}\|/\tau_{s})}{\sum_{k=1}^{|\mathcal{P}|}\exp(-\|l^{\prime}_{i}-g_{p_{k}}\|/\tau_{s})}$ (5)
where $\tau$ is a temperature parameter that controls the distribution’s sharpness and $g_{p_{j}}$ denotes the geographic coordinate of POI $p_{j}$. Concurrently, the temporal attention captures time-specific visit patterns. The temporal embedding $t^{\prime}_{vec,i}$ is projected into a query vector $w_{i}=W_{t}t^{\prime}_{vec,i}$, and the attention weight is computed via dot-product similarity:
$\alpha_{i,j}^{(t)}=\frac{\exp(f_{j}\cdot w_{i})}{\sum_{k=1}^{|\mathcal{P}|}\exp(f_{k}\cdot w_{i})}$ (6)
Finally, we adaptively balance the spatial and temporal signals by computing fusion weights using a softmax over their average representations $(\beta_{s},\beta_{t})=\text{Softmax}\left(\mathbf{W}_{f}\left[\frac{1}{L}\sum_{i=1}^{L}\mathbf{l}_{i}\,\|\,\frac{1}{L}\sum_{i=1}^{L}\mathbf{t}_{i}\right]\right)$ , where $\mathbf{l}_{i}$ and $\mathbf{t}_{i}$ denote the $i$-th spatial and temporal feature vectors, respectively; $\mathbf{W}_{f}$ is a learnable linear projection matrix that is used to create a unified attention distribution $\boldsymbol{\alpha}_{i,j}=\beta_{s}\boldsymbol{\alpha}^{(s)}_{i,j}+\beta_{t}\boldsymbol{\alpha}^{(t)}_{i,j}$. This fused attention is applied to the shared POI embeddings to create a context-aware representation $h_{\text{poi},i}=\sum_{j=1}^{|\mathcal{P}|}\alpha_{i,j}e_{\text{POI}}^{(j)}$. This POI-enhanced vector $h_{poi}$ is integrated with the initial features $h_{vec}=[l_{vec},t_{vec}]$ through a residual connection, and is processed by a Transformer Encoder to produce the contextualized representations $H$.
Table 1: Fidelity Evaluation. The best results on each dataset are in bold, and the second-best results are underlined.
Method Metric Distance Radius Interval Length Distance Radius Interval Length
FS-NYC FS-TKY
SMM (Maglaras and Katsaros 2015) 0.196 0.259 0.381 0.358 0.178 0.277 0.217 0.329
TimeGEO (Jiang et al. 2016) 0.251 0.625 0.278 0.386 0.277 0.524 0.232 0.216
Hawkes (Laub, Taimre, and Pollett 2015) 0.190 0.517 0.339 0.253 0.157 0.315 0.269 0.281
LSTM (Rossi et al. 2021) 0.166 0.441 0.175 0.235 0.184 0.292 0.230 0.257
SeqGAN (Yu et al. 2017) 0.116 0.364 0.150 0.291 0.108 0.292 0.207 0.276
MoveSim (Feng et al. 2020) 0.080 0.289 0.133 0.391 0.079 0.256 0.114 0.353
DiffTraj (Zhu et al. 2023b) 0.051 0.085 0.104 0.320 0.276 0.201 0.116 0.744
ControlTraj (Zhu et al. 2024) 0.193 0.356 0.160 0.337 0.049 0.096 0.191 0.223
Our Method (ours) 0.079 0.065 0.067 0.127 0.015 0.043 0.081 0.120
FS-ATX GW-STO
SMM (Maglaras and Katsaros 2015) 0.603 0.625 0.139 0.347 0.637 0.599 0.697 0.428
TimeGEO (Jiang et al. 2016) 0.416 0.501 0.148 0.401 0.579 0.489 0.727 0.397
Hawkes (Laub, Taimre, and Pollett 2015) 0.517 0.556 0.261 0.521 0.621 0.496 0.579 0.433
LSTM (Rossi et al. 2021) 0.589 0.602 0.187 0.339 0.618 0.527 0.699 0.369
SeqGAN (Yu et al. 2017) 0.452 0.475 0.029 0.292 0.531 0.462 0.745 0.413
MoveSim (Feng et al. 2020) 0.306 0.382 0.078 0.275 0.181 0.265 0.318 0.138
DiffTraj (Zhu et al. 2023b) 0.061 0.104 0.241 0.702 0.114 0.275 0.802 0.306
ControlTraj (Zhu et al. 2024) 0.108 0.217 0.150 0.286 0.204 0.253 0.765 0.141
Our Method (ours) 0.058 0.088 0.036 0.200 0.136 0.197 0.292 0.080
Multi-task Hybrid-Head Decoder
The Multi-task Hybrid-Head Decoder autoregressively generates the fine-grained trajectory, i.e., a sequence of (POI, timestamp) pairs, conditioned on the contextual representations $H$ from the encoder. To generate the $n^{\prime}$-th check-in point, the previously generated sequence $s^{\prime}_{\mathit{syn}}=[(p_{1},t_{1}),\ldots,(p_{n^{\prime}-1},t_{n^{\prime}-1})]$ is first embedded into a sequence of vectors $s^{\prime}_{vec}$, where each POI is replaced by its shared embedding and each timestamp is converted using Time2Vec. Next, a Transformer-based Decoder processes this embedded sequence and the encoder’s output $H$ to produce a hidden state $h_{n^{\prime}}$, which summarizes the necessary history and context. This hidden state is then fed into two specialized prediction heads to determine where and when the next check-in occurs. For POI prediction, a linear layer computes logits over the vocabulary: $\text{logits}_{\text{POI}}=W_{\text{POI}}h_{n^{\prime}}+b_{\text{POI}}$. For timestamp prediction, we model the temporally irregular nature of check-ins using a Neural Temporal Point Process (NTPP)  (Shchur et al. 2021). Unlike direct regression, which provides a single, often unrealistic point estimate by simply predicting an average time interval. The TPP models the full probability distribution of the next event by predicting the conditional intensity function $\lambda(t)=\text{Softplus}(W_{\text{time}}h_{n^{\prime}}+b_{\text{time}})$, which defines the instantaneous event rate given the current context. To sample the next timestamp, we first derive the cumulative distribution function $F^{-1}$ corresponding to $\lambda(t)$, and then draw a sample from this distribution via inverse transform sampling:
$t_{n^{\prime}}=F^{-1}(u),\quad\text{where }u\sim\mathcal{U}(0,1).$ (7)
This enables the model to generate diverse and realistic non-uniform time intervals, capturing the stochastic nature of real-world check-in behaviors.
Training Objective
Coarse2FineNet is trained with teacher forcing and causal masking (Vaswani et al. 2017). Our training objective is formulated as:
$\mathcal{L}=\lambda_{\text{POI}}\mathcal{L}_{\text{POI}}+\lambda_{\text{time}}\mathcal{L}_{\text{time}}+\lambda_{\text{spatial}}\mathcal{L}_{\text{spatial}}.$ (8)
The POI prediction loss $\mathcal{L}_{\text{POI}}=-\sum_{t=1}^{T}y_{t}\log(\hat{y}_{t})$ enforces categorical accuracy through cross-entropy, where $T$ represents the number of check-ins in a trajectory and $y_{t}$, $\hat{y}_{t}$ denote ground-truth and predicted POI id respectively. The temporal loss is the negative log-likelihood objective for the neural temporal point processes, formulated as $\mathcal{L}_{\text{time}}=-\sum_{i=1}^{T}(\log\lambda_{i}-\lambda_{i}\Delta t_{i})$ , where $\lambda_{i}$ is the intensity predicted by the network for the $i$-th interval, and $\Delta t_{i}$ is the ground-truth time duration since the last event. This loss maximizes the probability of the observed event sequence by training the model to predict a high intensity just before an event occurs. The spatial consistency loss $\mathcal{L}_{\text{spatial}}=\frac{1}{T}\sum_{t=1}^{T}\|p_{t}-p_{\hat{t}}\|_{2}$ penalizes geographic implausibility by measuring Euclidean distance between predicted and ground-truth POI locations, where $p_{t}$ and $p_{\hat{t}}$ denote the geographic coordinates of the ground-truth and predicted POIs, respectively. The three components are balanced by three learnable weights, enabling the model to generate trajectories that are spatially plausible, temporally accurate, and categorically consistent.
Evaluation
(a) CDF of distance
(b) CDF of radius
(c) CDF of length
(d) CDF of interval
Figure 4: Performance comparison of different methods on FS-NYC dataset.
Experiment Setup
Datasets
We utilize four publicly available LBSN datasets containing POI check-in trajectories to evaluate the performance of our proposed GeoGen, including Foursquare New York City (FS-NYC), Foursquare Tokyo (FS-TKY), Foursquare Austin (FS-ATX) (Yang et al. 2014; Yang, Zhang, and Qu 2016), and Gowalla (GW-STO) (Cho, Myers, and Leskovec 2011). We first truncate each individual LBSN check-in trajectory with a pre-defined duration (one week for the three FourSquare datasets and six weeks for GW-STO) and maintain the sequences with more than 10 visits to ensure data quality. Detailed data descriptions are shown in the Appendix. We split each dataset into training, validation, and testing sets with a ratio of 7:2:1.
Baselines
We compare our GeoGen with the following eight state-of-the-art models: SMM (Maglaras and Katsaros 2015), TimeGEO (Jiang et al. 2016), Hawkes (Laub, Taimre, and Pollett 2015), LSTM (Rossi et al. 2021), SeqGAN (Yu et al. 2017), MoveSim (Feng et al. 2020), DiffTraj (Zhu et al. 2023b), ControlTraj (Zhu et al. 2024). Detailed descriptions of these baselines are shown in the Appendix.
Metrics
We evaluate the synthetic data quality from both fidelity and utility perspectives. Following common practice (Zhu et al. 2023b; Feng et al. 2020), we evaluate data fidelity by computing the Jensen-Shannon Divergence (JSD) (Lin 1991) between the distributions of synthetic and real data across four metrics from both spatial and temporal aspects: 1) Distance: the distance between consecutive check-ins; 2) Radius: the radius of gyration, which measures the trajectory’s spatial range; 3) Interval: the time interval between consecutive check-ins; 4) Length: the total number of points in the trajectory. A lower JSD score indicates a higher fidelity of the synthetic data. We evaluate the utility of synthetic data by conducting data-driven downstream applications.
Overall Performance
Fidelity Evaluation
The comprehensive quantitative results are presented in Table 1. We found that traditional rule-based methods (SMM, TimeGEO, Hawkes) generally fall short, as their reliance on fixed statistical assumptions prevents them from capturing the complex, long-term behaviors from POI check-in trajectories. Designed for spatially continuous and temporally regular trajectory data generation, DiffTraj shows competitive performance on spatial metrics on dense urban datasets like FS-NYC, but performs poorly on the length metric due to the lack of awareness of discrete POI locations. Although the ControlTraj can model the connection between POIs as road segments, its incompatibility with temporally irregular data degrades the LBSN check-in trajectory generation quality. In contrast, our proposed model, GeoGen, significantly outperforms baselines across the four diverse datasets, highlighting its robustness and generalizability to generate high-fidelity LBSN trajectory data. For instance, on the FS-TKY dataset, GeoGen improves the second-best method by over 69% (0.049 - 0.015 / 0.049) in the Distance metric and 55% in the Radius metric. Similarly, it achieves a 38% improvement in trajectory Length on the GW-STO dataset.
The Cumulative Distribution Function (CDF) curves in Figure 4 (and more results in the Appendix) show that GeoGen’s distributions most closely align with those of the original data, confirming its effectiveness in capturing realistic check-in behavior patterns.
Table 2: Synthetic data utility for next check-in prediction.
Datasets FS-NYC FS-TKY
Methods Metrics RMSE ED RMSE ED
Real 0.187 7.01 0.235 13.70
SeqGAN 0.426 16.34 0.689 23.12
MoveSim 0.371 10.28 0.671 16.90
DiffTraj 0.424 15.27 0.728 18.28
ControlTraj 0.321 14.29 0.491 16.28
GeoGen 0.225 9.21 0.352 15.82
Utility Evaluation
As shown in Table 2, we evaluate the utility of the generated synthetic data through a next check-in prediction task (Yuan et al. 2023), where a prediction model is trained on synthetic data and tested on real-world data. Root Mean Square Error (RMSE) and Euclidean Distance (ED) are used to measure the prediction performance from the temporal dimension and spatial dimension, respectively. Our results demonstrate that GeoGen achieves the closest performance to the real data across both temporal and spatial metrics using FS-NYC and FS-TKY datasets, indicating its strong ability to preserve data utility. MoveSim performs better than SeqGAN and ControlTraj in the spatial dimension due to its explicit consideration of spatial structures such as distances between locations, while SeqGAN and ControlTraj treat locations as discrete categorical variables. Our proposed GeoGen consistently achieves the best performance across all metrics and datasets, thanks to its ability to effectively model spatially discrete and temporally irregular patterns of LBSN check-in trajectories.
In-depth Analysis
Ablation Study
To validate our S2TDiff in stage 1, we benchmark it against two other strategies: a) replication (S2TDiff-Rep): This method fills empty time slots with the coordinates of the last known check-in. b) fixed point (S2TDiff-Fix): This method assigns a predetermined central coordinate to all empty time slots. We also build a variant S2TDiff w/o S2G where the core module S2G Attention module is removed to test its effectiveness. For Coarse2FineNet in the second stage, we build a variant C2F w/o DCF by removing the dynamic context fusion component. The results on the FS-NYC dataset are shown in Table 3. We found that interpolation can provide richer spatio-temporal patterns and underlying physical constraints compared with replication and fix point latent movement sequence reconstruction strategies. Meanwhile, the performance decreased in both spatial and temporal domains without the S2G Attention module, also proving its effectiveness compared with pure convolutions in the U-Net. In Coarse2FineNet, removing the dynamic context fusion and POI embedding (C2F w/o DCF) causes a dramatic performance decrease due to the lack of informative supervision for the learning of the discrete POI embeddings.
Table 3: Performance comparison across different GeoGen variants on FS-NYC dataset.
Methods Metrics Distance Radius Interval Length
$\text{S}^{2}\text{TDiff-Rep}$ 0.084 0.071 0.074 0.128
$\text{S}^{2}\text{TDiff-Fix}$ 0.097 0.073 0.079 0.130
$\text{S}^{2}\text{TDiff w/o S}^{2}\text{G}$ 0.087 0.069 0.080 0.135
C2F w/o DCF 0.119 0.138 0.096 0.169
Complete GeoGen 0.079 0.065 0.067 0.127
Impact of Granularity on Model Performance and Efficiency
We investigate the impact of time intervals $I$. As shown in Table 4, we found that smaller intervals (e.g., 2-4 hours) yield better temporal fidelity (Interval), whereas spatial fidelity metrics (Distance and Radius) are optimized at a larger 6-hour interval. Increasing the interval from 2 to 6 hours reduces Stage 1 memory consumption from $2.36$ GB to $0.35$ GB and increases throughput by more than fourfold (from $11,322$ to $50,540$ samples/sec). In the second stage, the computational cost of Coarse2FineNet remains relatively stable across different intervals.
Table 4: Results under different coarse-grained interval II selection on GW-STO dataset with batch size 256 for S2TDiff (S1) and 16 for Coarse2FineNet (S2). The unit of throughput is samples per second.
Metric I 2 hours 3 hours 4 hours 6 hours 8 hours 12 hours
Distance 0.454 0.183 0.136 0.073 0.089 0.080
Radius 0.402 0.250 0.197 0.154 0.153 0.154
Interval 0.292 0.292 0.292 0.315 0.342 0.444
Length 0.318 0.137 0.080 0.115 0.134 0.131
S1 Memory (GB) 2.36 1.14 0.69 0.35 0.21 0.12
S2 Memory (GB) 17.55 9.78 10.10 10.68 3.62 3.92
S1 Throughput 11322 20221 30058 50540 73794 119932
S2 Throughput 39 41 44 45 49 49
Related Work
LBSN Data Mining
In recent years, LBSN data have attracted much interest from both research and industry communities (Ye, Yin, and Lee 2010; Ding et al. 2018) due to their high value for many practical applications such as POI recommendations (Luo, Liu, and Liu 2021), user behavior analysis (Ding et al. 2018), and advertising (Jeon et al. 2021). Yang et al. (Yang et al. 2024) design a self-explainable algorithm to recommend the next visiting POI with high accuracy based on LBSN data. Han et al. (Han et al. 2021) leverage LBSN data to understand human activities during COVID-19 for better pandemic control strategies. However, given the potential privacy concerns and high collection costs, it becomes highly challenging for researchers to access large-scale LBSN data (Long et al. 2024; Cai et al. 2024). Hence, in this study, we focus on generating large-scale synthetic LBSN data that shares similar data characteristics with real data and also has high utility for downstream applications.
Trajectory Data Generation
The recent advances in generative models provide a great opportunity for synthetic trajectory data generation. These models (e.g., Generative Adversarial Network (GAN) (Rao et al. 2020; Liu, Chen, and Andris 2018; Cao and Li 2021) and diffusion models (Zhu et al. 2023b)) utilize the power of deep neural networks to learn and replicate the underlying spatio-temporal distributions of real trajectory data. Once trained, these models generate synthetic trajectories by sampling from the learned distributions. For example, TrajGen (Cao and Li 2021) utilizes GAN and a Seq2Seq model to generate spatial-temporal trajectory data. Feng et al. (Feng et al. 2020) propose a GAN-based framework, MoveSim, with the introduction of prior knowledge of human mobility. MoveSim (Feng et al. 2020) generates the next location based on the last point in the partially generated sequence and iteratively generates the rest trajectory points. Zhu et al. (Zhu et al. 2023b, a, 2024) propose a pioneering work by adopting the diffusion model to generate continuous and road network-constrained trajectories with a fixed length and time interval between two trajectory points. Different from existing works that usually focus on generating trajectories in continuous spatial spaces with fixed time intervals and fixed lengths, our work aims to generate spatially discrete, temporally irregular LBSN check-in trajectories.
Conclusion
In this paper, we propose GeoGen, a novel two-stage coarse-to-fine framework for the generation of synthetic LBSN trajectories that are characterized by spatial discreteness and temporal irregularity. In the first stage, we propose a Sparsity-aware Spatio-temporal Diffusion model (S2TDiff) with an efficient Spatially-Aware Sparsely-Gated U-Net (SASG-UNet) to learn multi-scale behavioral patterns. In the second stage, we design Coarse2FineNet, a Transformer-based Seq2Seq architecture equipped with a Dynamic Context Fusion mechanism in the encoder and a multi-task hybrid-head decoder to generate accurate next POI location and its fine-grained timestamp. We extensively evaluate our GeoGen with four datasets, and results show that GeoGen outperforms state-of-the-art models from both data fidelity and utility perspectives, e.g., it increases 69% and 55% on distance and radius metrics on the FS-TKY dataset, and enhances the next check-in prediction task.
Appendix
Datasets
Table 5 presents detailed statistics for the four datasets used in our study. To ensure consistency and mitigate the impact of latent mobility pattern changes over extended periods, we filtered trajectories from specific time ranges: April 2012 to June 2012 for the Foursquare datasets (FS-NYC, FS-TKY, FS-ATX) and February 2009 to October 2010 for the Gowalla Stockholm dataset (GW-STO). Additionally, we excluded trajectories shorter than 10 check-ins for FS-NYC, FS-TKY, and GW-STO, and shorter than five check-ins for FS-ATX.
Table 5: Dataset statistics. Traj is the abbreviation of trajectory in the table.
FS-NYC FS-TKY FS-ATX GW-STO
# of POIs 18207 24689 8591 11803
# of Check-ins 78334 144999 71382 63211
# of Trajs 4482 7576 319 2094
Duration 1 week 1 week 1 week 6 weeks
Mean Traj Length 17 19 12 30
Unique POIs / Traj 11 12 9 20
Motivation and Data Analysis
Notably, many existing studies (Yu et al. 2017; Feng et al. 2020; Zhu et al. 2023b; Feng et al. 2020; Zhu et al. 2024) rely on temporally regular trajectories to generate temporally irregular LBSN check-in trajectories through linear interpolation. As shown in Figure 5, we conduct data analysis to explore the statistical differences between original LBSN check-in trajectories and the interpolated counterparts with different time intervals between consecutive trajectory points. The processed trajectories with a one-minute interval can be regarded as fine-grained LBSN check-in trajectories, while others are coarse-grained. Though the interpolation has a relatively smaller decrease in spatial metrics, it has more influence on length and interval metrics. For example, with the increase of the interval, more check-ins occurring within the interval are overlapped. As depicted in Figure 5(d), the interval distribution of the reconstructed trajectories differs significantly from the original: the absence of short intervals and the unsmooth curve both contribute to the loss of useful temporal information. Similarly, Figure 5(c) clearly illustrates the missing points, leading to a loss of spatio-temporal information. Thus, the interpolated LBSN check-in trajectories significantly influence the quality of the data due to the missing points. However, on the other hand, it is challenging for these methods to generate long-term, fine-grained LBSN check-in trajectories with a short interval due to the high dimensionality of long-term synchronous trajectory data. Therefore, a possible solution is to reduce the data’s dimensionality by increasing the interval, but it inevitably leads to a coarser granularity in the generated LBSN check-in trajectories.
(a) CDF of distance
(b) CDF of radius
(c) CDF of length
(d) CDF of interval
Figure 5: CDF curves of interpolated LBSN check-in trajectories with different intervals.
Baselines
We utilize the following eight methods as baselines for the evaluation of our proposed GeoGen:
  • SMM (Maglaras and Katsaros 2015): Semi-Markov process models the interval between two adjacent visits by the exponential distribution. The spatial domain is modeled by the transition matrix and the intensity of the interval.
  • TimeGEO (Jiang et al. 2016): TimeGEO captures the temporal domain by home-based context information and the spatial domain by the explore and preferential return (EPR) model.
  • Hawkes (Laub, Taimre, and Pollett 2015): Hawkes process is a temporal point process model. The partially observed sequence will influence the intensity function of the points to be predicted.
  • LSTM (Rossi et al. 2021): It generates a trajectory by predicting the time interval and location of visits sequentially.
  • SeqGAN (Yu et al. 2017): SeqGAN is a state-of-the-art GAN-based discrete sequence generation model.
  • MoveSim (Feng et al. 2020): MoveSim combines the advantages of model-based and model-free methods by introducing prior knowledge of human mobility regularities into a GAN-based framework.
  • DiffTraj (Zhu et al. 2023b): DiffTraj is a diffusion-based model for generating synchronous GPS trajectories.
  • ControlTraj (Zhu et al. 2024): ControlTraj is a diffusion-based framework that generates high-fidelity, controllable trajectories by integrating structural constraints from road network topology.
SMM, TimeGEO, and Hawkes are physics-driven approaches that incorporate physical rules to model spatio-temporal data. In contrast, the remaining methods are learning-based, employing deep neural networks to learn the distribution of POI check-in trajectories.
Metrics
We evaluate the quality of the synthetic trajectories from two perspectives: fidelity, which measures how realistically the synthetic data replicates real-world patterns; and utility, which assesses the usefulness of the generated data for a downstream prediction task.
Fidelity. To assess fidelity, we compare key spatial and temporal distributions between the real and synthetic datasets using the Jensen-Shannon Divergence (JSD). JSD is a symmetric and smoothed version of the Kullback-Leibler (KL) divergence, designed to quantify the similarity between two probability distributions. Specifically, the JSD between distributions $P$ and $Q$ is defined as:
$\mathrm{JSD}(P||Q)=\frac{1}{2}D_{\mathrm{KL}}(P||M)+\frac{1}{2}D_{\mathrm{KL}}(Q||M),$
where $M=\frac{1}{2}(P+Q)$ is the average distribution, and $D_{\mathrm{KL}}(P||Q)$ is the Kullback-Leibler divergence:
$D_{\mathrm{KL}}(P||Q)=\sum_{i}P(i)\log\frac{P(i)}{Q(i)}.$
JSD is always bounded between 0 and 1 (for distributions using log base 2), and a lower JSD value indicates that the synthetic data more closely resembles the statistical properties of the real data in terms of spatial and temporal patterns.
Following common practice (Zhu et al. 2023b; Feng et al. 2020), we compute the JSD over four spatio-temporal metrics that capture both local and global patterns in LBSN check-in trajectories. Given a check-in trajectory $s=[x_{1},x_{2},\ldots,x_{n}]$, where each check-in point $x_{i}=(p_{i},t_{i})$ consists of a POI ID $p_{i}$ associated with a geographic coordinate $\mathbf{g}_{p_{i}}$, and a timestamp $t_{i}$, we define the four evaluation metrics as follows:
  • Distance: the total travel distance of a trajectory, computed as
    $D=\sum_{i=2}^{n}\|\mathbf{g}_{p_{i}}-\mathbf{g}_{p_{i-1}}\|_{2},$
    which captures the cumulative spatial movement across all consecutive check-ins.
  • Radius: half of the maximum pairwise distance between any two check-in locations in the trajectory, i.e.,
    $R=\frac{1}{2}\max_{1\leq i<j\leq n}\|\mathbf{g}_{p_{i}}-\mathbf{g}_{p_{j}}\|_{2},$
    representing the spatial extent or spread of the trajectory.
  • Interval: the sequence of time gaps between consecutive check-ins, defined as $\Delta t_{i}=t_{i}-t_{i-1}$ for $i=2,\ldots,n$, reflecting the user’s temporal behavior.
  • Length: the number of check-in points in the trajectory, denoted as $\ell=n$, indicating the total trajectory length.
Utility. For utility evaluation, we measure performance on a next check-in prediction task. The spatial accuracy of predictions is assessed using the Euclidean Distance (ED), defined as
$ED(p,q)=\sqrt{(p_{x}-q_{x})^{2}+(p_{y}-q_{y})^{2}},$
where $p=(p_{x},p_{y})$ and $q=(q_{x},q_{y})$ denote the predicted and actual geographic coordinates of the next check-in, respectively. This metric provides an intuitive spatial error measure, where smaller distances indicate higher predictive precision.
In addition, we evaluate the accuracy of predicted timestamps using the Root Mean Square Error (RMSE), defined as
$RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$
where $y_{i}$ and $\hat{y}_{i}$ represent the ground truth and predicted timestamps for the $i$-th check-in, respectively, and $n$ is the total number of predicted check-ins. RMSE penalizes larger errors more heavily and reflects the average deviation between predicted and actual time values.
Across all three metrics, lower values indicate better performance, meaning the synthetic trajectories are both more realistic and more effective for downstream applications.
Implementation Details
We implemented the GeoGen framework in PyTorch and conducted all experiments on one NVIDIA A10 GPU with 24 GB memory. The model employs a two-stage training process to generate realistic human mobility trajectories.
In the first stage, we train the SASG-UNet model in S2TDiff for coarse-grained latent movement sequence generation using diffusion processes. The latent movement sequence is structured into tensors of shape $\mathbb{R}^{L\times 3}$, $L$ is 168 for the three FourSquare datasets and 1008 for the GW-STO dataset. The SASG-UNet architecture features an input/output channel dimension of 3, base channel dimension of 128, and channel multipliers of [1, 2, 2, 2] across resolution levels with 2 residual blocks per level and attention mechanisms at 16×16 resolution. The model uses exponential moving average (EMA) with a rate of 0.9 for stable training. This stage runs for 500 epochs using $T=1000$ diffusion timesteps with a linear noise schedule where $\beta_{1}=10^{-4}$ and $\beta_{T}=0.02$. We employ the AdamW optimizer with a learning rate of $\alpha=2\times 10^{-4}$. Training uses a batch size of 32 with validation performed every 10 epochs and model checkpoints saved every 50 epochs.
In the second stage, we train the Coarse2FineNet for fine-grained LBSN check-in trajectory generation. The model features a Time2Vec temporal embedding with 16 dimensions, encoder and decoder hidden dimensions of 64, 4 Transformer layers with 4 attention heads each, and 2 decoder layers. The architecture processes coarse trajectories with a maximum sequence length of 75 and supports up to 6 events per 60-minute interval. Spatial features utilize soft assignment with temperature parameter 0.25, while temporal features employ 1-minute base units with a minimum 1-minute interval between events. This stage trains for 100 epochs using the Adam optimizer with a learning rate of $\alpha=10^{-3}$, gradient clipping at 1.0, and learning rate reduction on plateau with factor 0.5 and patience of 5 epochs.
For our baseline comparisons, we report the key hyperparameters selected to ensure fair and competitive performance. The SMM (Maglaras and Katsaros 2015) model’s number of states was set to the total number of unique POIs in each respective dataset. For TimeGEO (Jiang et al. 2016), following the original author’s suggestions, we set the return probability $\rho=0.6$ and the exploration probability $\gamma=0.2$. The parameters for the Hawkes (Laub, Taimre, and Pollett 2015) process, which utilizes an exponential kernel, were optimized via maximum likelihood estimation on the training data. The deep learning methods, LSTM (Rossi et al. 2021), SeqGAN (Yu et al. 2017), and MoveSim (Feng et al. 2020), were configured with a hidden dimension and embedding size of 128 and trained with a batch size of 64 using the Adam optimizer with a learning rate of $1\times 10^{-3}$. For SeqGAN, the number of Monte Carlo search rollouts was set to 16. To adapt DiffTraj (Zhu et al. 2023b) and ControlTraj (Zhu et al. 2024), originally designed for spatially continuous, temporally regular trajectory generation, for fair comparison, we follow the same latent movement sequence reconstruction process and extract the two spatial dimensions as the training and target data for them. They utilize $T=1000$ diffusion steps and a linear noise schedule from $\beta_{start}=10^{-4}$ to $\beta_{end}=0.02$. For ControlTraj, we set the constraint guidance scale to 1.5. The backbone network for both diffusion baselines is a standard UNet architecture.
Extended Results
In this section, we provide additional visualizations of the generated POI check-in trajectories.
Figure 6 presents CDF visualizations of spatial and temporal metrics for FS-TKY, FS-ATX, and GW-STO datasets. We found that GeoGen consistently outperforms all baseline methods across all metrics. Due to the similar scale and characteristics of FS-TKY and FS-NYC, the CDF curves for FS-TKY follow a comparable trend, though with a slightly shorter travel radius. In contrast, FS-ATX presents a greater challenge due to its limited size, e.g., only 319 trajectories and 8,591 unique POIs. In this low-resource setting, GeoGen achieves the most balanced performance across all four spatio-temporal metrics. Notably, it closely replicates the distributions of the two spatial metrics (distance and radius), while other methods suffer noticeable performance drops. The GW-STO dataset contains longer trajectories (averaging 20 check-ins), resulting in more complex spatio-temporal patterns. While DiffTraj slightly outperforms GeoGen on the distance metric (0.114 vs. 0.136), GeoGen leads on the remaining three metrics, highlighting its robust modeling capabilities. Overall, GeoGen delivers superior performance across diverse datasets and metrics, demonstrating its generalizability and effectiveness.
To further validate the generalizability of GeoGen, we conduct the same ablation study on the GW-STO dataset. The results are summarized in Table 6. We observe consistent trends with those reported in the main content of the paper. The full GeoGen model outperforms all ablated variants across all evaluation metrics, demonstrating its robustness across datasets. Specifically, both the replication and fixed-point strategies lead to significantly inferior performance in capturing meaningful spatio-temporal dynamics, as reflected by higher errors in Distance, Radius, and Interval. Moreover, removing the S2G Attention module results in increased reconstruction errors, especially in the Interval metric, indicating that the module is crucial for modeling fine-grained temporal dependencies. Similarly, removing the dynamic context fusion component in Coarse2FineNet (C2F w/o DCF) leads to the most substantial performance drop, confirming the importance of contextual supervision in learning POI-aware fine-grained transitions.
Table 6: Performance comparison across different GeoGen variants on GW-STO dataset.
Methods Metrics Distance Radius Interval Length
$\text{S}^{2}\text{TDiff-Rep}$ 0.141 0.221 0.318 0.099
$\text{S}^{2}\text{TDiff-Fix}$ 0.159 0.216 0.320 0.097
$\text{S}^{2}\text{TDiff w/o S}^{2}\text{G}$ 0.162 0.220 0.352 0.106
C2F w/o DCF 0.213 0.256 0.338 0.123
Complete GeoGen 0.136 0.197 0.292 0.080
Table 7: Results under different coarse-grained interval II selection on FS-NYC dataset with batch size 256 for S2TDiff (S1) and 16 for Coarse2FineNet (S2). The unit of throughput is samples per second.
Metric I 30 mintues 1 hour 2 hours 3 hours 4 hours
Distance 0.082 0.079 0.078 0.088 0.092
Radius 0.081 0.065 0.071 0.073 0.153
Interval 0.063 0.067 0.092 0.091 0.115
Length 0.139 0.127 0.137 0.141 0.143
S1 Memory (GB) 1.19 0.32 0.13 0.06 0.05
S2 Memory (GB) 9.15 10.18 3.81 2.48 1.12
S1 Throughput 21009 50156 121034 186291 252140
S2 Throughput 41 45 50 52 57
We provide extended results on the impact of different coarse-grained time intervals $I$ using the FS-NYC dataset, as summarized in Table 7. The results show that temporal fidelity (measured by the Interval metric) is highest at finer intervals, with the best score (0.063) at 30 minutes and strong performance at 1 hour (0.067). In contrast, spatial fidelity metrics such as Distance and Radius achieve their lowest errors at intermediate intervals, particularly at 1 hour (Radius = 0.065) and 2 hours (Distance = 0.078), indicating a favorable trade-off between spatial accuracy and temporal resolution. Regarding computational efficiency, S2TDiff benefits from increasing interval length, with memory consumption dropping from 1.19 GB at 30 minutes to just 0.05 GB at 4 hours. Correspondingly, throughput improves substantially from 21,009 to 252,140 samples per second. Coarse2FineNet shows relatively stable resource use and throughput across intervals, reflecting its reliance on fixed coarse check-in counts rather than interval granularity.
Figure 7 compares the POI check-in distributions from real LBSN data against those from synthetic trajectories generated by various methods. Clearly, the check-in patterns from our proposed GeoGen model most closely match the real-world distributions across all four datasets. In contrast, ControlTraj fails to simulate the POI visiting distribution, as it lacks a specific mechanism for modeling points of interest. While MoveSim captures the general shape of the distributions, it cannot reproduce the high check-in density of the most popular POIs in the FS-NYC, FS-TKY, and FS-ATS datasets.
(a) CDF of distance in FS-TKY
(b) CDF of radius in FS-TKY
(c) CDF of length in FS-TKY
(d) CDF of interval in FS-TKY
(e) CDF of distance in FS-ATX
(f) CDF of radius in FS-ATX
(g) CDF of length in FS-ATX
(h) CDF of interval in FS-ATX
(i) CDF of distance in GW-STO
(j) CDF of radius in GW-STO
(k) CDF of length in GW-STO
(l) CDF of interval in GW-STO
Figure 6: Performance comparison on the other three datasets.
(a) Density distribution on FS-NYC dataset.
(b) Density distribution on FS-TKY dataset.
(c) Density distribution on FS-ATX dataset.
(d) Density distribution on GW-STO dataset.
Figure 7: LBSN POI check-in distributions of different methods on four datasets. The brighter the color is, the more frequently the POI is visited.